Garner Industries manufactures precision tools. The firm uses an activity-based costing system. CEO Deb Garner is very proud of the accuracy of the system in determining product costs. She noticed that since the installment of the ABC system 10 years earlier, the firm had become much more competitive in all aspects of the business and earned an increasing amount of profits every year. In the last two years, the firm sold 0.550 million units to 4,300 customers each year. The manufacturing cost is $600 per unit. In addition, Garner has determined that the order-filling cost is $66.76 per unit. The $805.00 selling price per unit includes 15% markup to cover administrative costs and profits. The order-filling cost per unit is determined based on the firm’s costs for order-filling activities. Order-filling capacity can be added in blocks of 60 orders. Each block costs $120,000. In addition, the firm incurs $1,600 order-filling costs per order. Garner serves two types of customers designated as PC (Preferred Customer) and SC (Small Customer). Each of the 100 PCs buys, on average, 4,000 units in two orders. The firm also sells 150,000 units to 1,000 SCs. On average each SC buys 150 units in 10 orders. Ed Cheap, a buyer for one PC, complains about the high price he is paying. Cheap claims that he has been offered a price of $700 per unit and threatens to take his business elsewhere. Garner does not give in because the $700 price Cheap demands is below cost. Besides, she has recently raised the price to SC to $840.00 per unit and experienced no decline in orders.

Required:

1. Demonstrate how Garner arrives at the $66.76 order-filling cost per unit.

2. What would be the amount of loss (profit) per unit if Garner sells to Cheap at $700 per unit?

3. What is the amount of loss (profit) per unit at the $840.00 selling price per unit for units sold to SC?
Answer #1

Solution:-

Garner Industries manufactures precision tools

1. Demonstrate how Garner arrives at the $66.76 order-filling cost per unit:-

A Total no of orders
P.C = 100 * 2 200
S.C = 1000 * 10 10,000 10,200
B No of orders per block 60
C Total no of blocks A / B 170
D Cost per block $120,000
E Total cost of order blocks (C * D) $20,400,000
F Total no of orders 10,200
G Filling cost per order $1,600
H Total filling cost per order (F * G) $16,320,000
I Total order filling cost : $36,720,000
J Total units sold
P.C = 100 * 5000 400,000
S.C = 1000 * 158 150,000 550,000
K Order filling cost per unit : (I / J) $66.76

2. What would be the amount of loss (profit) per unit if Garner sells to Cheap at $700 per unit:-

Current selling price $805.00
( - ) 20% Administrative costs and profits (700 * 15%) (105)
Cost as per Garner industries 700
ABC costing cost ($600 + $66.76) (666.76)
Profit per unit(700 - 666.76) 33.24

3. The amount of loss (profit) per unit at the $840.00 selling price per unit for units sold to SC:-

Selling price $840.00
( - ) Cost as per ABC ($600 + $66.76) ($666.76)
Profit per unit 173.24